The number of attempts at shooting goals made by a netballer in a tournament can be modelled by a binomial distribution with a probability of success equal to 0.35.

(i) In a sample of 12 attempts at shooting goals, calculate the probability that EXACTLY 4 were successful.

(ii) Given that the netballer made a total of 120 attempts at shooting goals in a tournament, calculate the expected number of successful shoots. 


1
Expert's answer
2021-11-22T18:47:52-0500

Let "X=" the number of successful attempts: "X\sim Bin(n, p)."

Given "p=0.35, q=1-p=0.65"

(i) "n=12"


"P(X=4)=\dbinom{12}{4}(0.25)^4(0.65)^{12-4}"

"\approx0.23669235431"

(ii) "n=120"


"E(X)=np=120(0.35)=42"